Karen is weighing different children’s blocks. Four red blocks and
five blue blocks weigh 32 ounces. One red block and eight blue
blocks weigh 35 ounces. What is the total weight of one red and
one blue block?

Found 2 solutions by chen, josgarithmetic:
Answer by chen(11)

4r+5b=32
r+8b=35 so r=35-8b. We substitute to the 1st:
4(35-8b)+5b=32 -->
140-32b+5b=32 --> 27b=108 --> b=4 so r=35-32=3
So r+b=7

Answer by josgarithmetic(39618)

Identical, already done:
https://www.algebra.com/algebra/homework/word/misc/Miscellaneous_Word_Problems.faq.question.1081519.html
https://www.algebra.com/algebra/homework/word/misc/Miscellaneous_Word_Problems.faq.question.1081519.html